OpenStudy (oryzae):

cosX/secX+tanX = 1-sinX Please help! Proving with Identities.

Nnesha (nnesha):

is it \[\large\rm \frac{ cosx }{ \sec x }+ \tan x =1-\sin(x)\] `OR` \[\large\rm \frac{ cosx }{ \sec x +tan x } =1-\sin(x)\]

OpenStudy (oryzae):

Oooh, sorry! I forgot the parentheses. It's the second version, with both secX and tanX in the denominator.

Nnesha (nnesha):

alright first i would multiply the fraction by the conjugate of the denominator or first you can rewrite sec and tan in terms of sin, cos

Nnesha (nnesha):

what's the definition of sec nd tan ??

Nnesha (nnesha):

how would you write sec and tan in terms of sin,cos ? do you know ? if not its okay :)